# GeoLayer: Towards Low-Latency and Cost-Efficient Geo-Distributed Graph Stores with Layered Graph

### v1

@@ V-B Replica Placement Built on Layered Graph
Theoretical analysis. In Theorem 1, we provide a theoretical guarantee for the existence of a non-trivial steady-state solution in the heat system defined by the DHD model.
@@
$\text{where}\quad\small\Delta\hat{C}^{(A)}_{xy}=R_{xy}\cdot\left(\lambda_{1}\sum_{j=1,BS_{j}\in\mathcal{C}}^{n}\sigma_{xyj}\right).$
(a) Pattern overlap
(b) Overlap-centric diffusion
To proactively guard against sudden surges in access frequency, we use the DHD model to identify and pre-cache these high-heat regions. First, the DC issues a global unobserved edge request to upper layers to acquire complete topological connectivity of externally cached regions. Concurrently, it initializes a high-heat vertex set $\mathcal{O}=\{v\in V\mid\mathcal{H}_{v}^{0}\geq\theta\}$, treating these vertices as external heat sources as defined in Equation (9). The system then iteratively updates heat states using Equation (10) until it reaches steady state (see Equation (12)). At equilibrium, we extract the high-heat subgraph $G^{H}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v\in V\mid\mathcal{H}_{v}^{*}\geq\theta\}$ and $\mathcal{E}=\{(u,v)\mid u,v\in\mathcal{V}\}$. This subgraph is subsequently pre-cached in the DC.
Theoretical analysis. We analyze the necessity of pre-caching high-heat regions. Given $\mathcal{V}=\{v\in V\mid\mathcal{H}_{v}^{*}\geq\theta\}$ and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. Partition $\mathcal{E}$ into observed edges $\widetilde{\mathcal{E}}$ and unobserved edges $\mathcal{E}\setminus\widetilde{\mathcal{E}}$. For each $(u,v)\in\widetilde{\mathcal{E}}$, define transition probability as
@@ 14           $\textnormal{{UpdateRoutingTable}}(v)$
Implementation. Algorithm 2 outlines the replica placement process. Initially, any external pattern $p$ retrieved at a given layer $L_{k}$ is included in the replication procedure if it satisfies $R_{pL_{k}}>W_{pL_{k}}$. The process begins with the layer-by-layer sinking of pattern replicas, guided by latency requirements (Line 2). Next, the data replication gain $C^{rep}_{p}$ is computed to determine whether the pattern should be fully replicated or decomposed in the next lower layer (Lines 22). If $C^{rep}_{py}<0$, the pattern is decomposed across the corresponding BSs (Line 2). For pattern decomposition, patterns are first partitioned into a set of disjoint regions based on their overlapping components (Line 2). The replication gain ${C}^{rep}_{\omega}$ of each region is then evaluated to determine whether the region should be replicated across $BS$s or enter a competitive allocation process (Line 2). For regions subject to competition, each $BS$ executes the DHD model (Line 2), and region ownership is determined based on the resulting heat distribution (Line 2). This process proceeds layer by layer until all data items reach the DCs at layer $L_{0}$ (Line 2). Finally, high-heat regions are pre-cached in the DCs at layer $L_{0}$ (Line 2).
VI Stepwise Layered Routing
Figure 5: Request routing under the online query mode.
Figure 6: Request routing under the offline query mode.
VII Experimental evaluation
VII-A Experimental Setup
Figure 7: Request latency.
Figure 8: Comparison of cost metrics.
Datasets. Our evaluation employs three distinct datasets: a synthetic social network generated using the LDBC-SNB benchmark with a scale factor (SF) of 10, and two real-world graph datasets, UK-2005 [58] and Twitter-2010[59]. Table IV provides a detailed summary of the statistics for each dataset.
TABLE IV: Dataset Description
Graph Vertices Edges Abbreviation
LDBC-SNB-SF10 [60] 36,423,181 231,371,359 SNB
UK-2005 [58] 39,459,925 936,364,282 UK
Twitter-2010 [59] 41,652,230 1,468,364,884 TW
[…]
VII-B Effectiveness Analysis in Online Mode
[…]
We also report the algorithm execution time, as shown in Figure LABEL:fig:exec_time. Random-3 and Top-3 exhibit lower computational complexity due to their simple heuristic strategies. ADP incurs higher overhead as it relies on hypergraph partitioning and multiple rounds of iteration. Although $\mathsf{GeoLayer}$ involves additional complexity due to its multi-layered design, it remains efficient by leveraging cluster-level parallelism, producing results within a reasonable execution time.
[…]
VII-C Performance Comparison in Offline Mode
We assess request routing effectiveness for graph analysis.
[…]
WAN Cost. We evaluate WAN usage during algorithm execution by measuring the actual volume of inter-DC data transfer. As shown in Figure LABEL:fig:engine_WAN, $\mathsf{GeoLayer}$ incurs the lowest WAN cost across all test conditions. Specifically, $\mathsf{GeoLayer}$ reduces WAN cost by an average of 48.1% (up to 72.0%) compared to RAGraph, 26.3% (up to 44.7%) compared to RAGraph+, and 36.5% (up to 68.2%) compared to GrapH. The reduction in WAN usage stems from $\mathsf{GeoLayer}$’s ability to select fewer execution sites while minimizing the introduced cross–DC communication. $\mathsf{GeoLayer}$ constructs $2$ execution sites for the $\mathsf{SNB}$ graph, and $3$ sites for both the $\mathsf{UK}$ and $\mathsf{TW}$ graphs.
Migration Ratio. We also report the ratio of migrated edges under each strategy, as shown in Figure LABEL:fig:migration_rate. $\mathsf{GeoLayer}$ achieves a migration ratio ranging from 34% to 42% across the three datasets. On the $\mathsf{SNB}$ graph, $\mathsf{GeoLayer}$ exhibits a relatively higher migration ratio, primarily due to the consolidation of data into only two execution sites to enable more efficient processing. All evaluated strategies consider network heterogeneity in geo-distributed environments and aim to improve graph analysis performance by reducing communication overhead through limited data movement.
VIII Related Works
Data Replication and Request Routing. A large body of work [18, 19, 20, 21, 22, 23, 24, 67, 68, 69, 70, 71, 72, 73] focuses on optimizing replica placement and request routing decisions in geo-distributed environments. These efforts address the problem from multiple perspectives, considering factors such as replica placement [22, 23, 67, 24], caching time [18], data migration [65, 66], QoS/QoE/SLA guarantees [19, 20, 69], consistency models [21, 70], security policies [71, 72], and load balancing [73, 68]. Several studies further incorporate data association into decision-making processes [25, 17, 26, 27, 28, 29, 74]. Volley [74] proposes an iterative optimization algorithm based on weighted spherical means to handle inter-data dependencies. Yu et al. [28, 17] model data association as a hypergraph partitioning problem to guide co-location decisions. Liu et al. [27] develop a replica placement strategy based on overlapping community detection to handle associated data. These approaches commonly abstract associated data items as edges and connect them with storage nodes to form a graph structure, thereby enabling graph-based modeling of placement decisions. However, these associations are driven by job-level relationships, whereas $\mathsf{GeoLayer}$ focuses on the intrinsic structural properties of the graph.
Geo-Distributed Graph Data Processing. Many existing efforts [75, 76, 47, 77, 78, 79, 80, 46] aim to optimize the performance of graph-structured data processing in geo-distributed settings. GeoGraph [75] leverages clustered DCs and a metagraph to support graph query processing. PGPregel [47] adopts sampling and combiner techniques to enable privacy-preserving graph processing. Hosseinalipour et al. [77] propose an energy-efficient graph job assignment method for geo-distributed cloud networks. These works focus on optimizing the performance of upstream computing systems, and are thus orthogonal to $\mathsf{GeoLayer}$’s design. In parallel, several works explore geo-distributed graph partitioning to reduce communication cost. RLCut [78] and DistRLCut [79] employ reinforcement learning to address the complexity of dynamic graph partitioning. GeoCut [80] proposes a geography-aware partitioning method that accounts for inter-data center cost constraints. These methods are designed around specific graph tasks and can be naturally built upon the data distribution profiles generated by $\mathsf{GeoLayer}$’s offline-mode routing.
IX Conclusion

### v3

@@ IX-A Convergence of DHD Model
In Theorem 1, we provide a theoretical guarantee for the existence of a non-trivial steady-state solution in the heat system defined by the DHD model.
@@
IX-B Guarantees of Pre-Caching
We analyze the necessity of pre-caching high-heat regions. Given $\mathcal{V}=\{v\in V\mid\mathcal{H}_{v}^{*}\geq\theta\}$ and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. Partition $\mathcal{E}$ into observed edges $\widetilde{\mathcal{E}}$ and unobserved edges $\mathcal{E}\setminus\widetilde{\mathcal{E}}$. For each $(u,v)\in\widetilde{\mathcal{E}}$, define transition probability as
@@ IX-B Guarantees of Pre-Caching
IX-C Complexity Discussion
This section explains how each component of $\mathsf{GeoLayer}$ acts in concert to shrink the effective complexity of solving the objective in Equation (6).
The joint BIP formulation over the variable set $\{\delta,\sigma,\rho\}$ couples replica placement and routing across both DCs and query patterns, which is theoretically NP-hard. The goal of $\mathsf{GeoLayer}$ is not to compute an exact global optimum, but rather to reduce effective problem complexity through structured decomposition:
(i) Latency-driven sinking shrinks candidates. Let $\mathcal{C}_{k}$ be the set of clusters at layer $L_{k}$. The sinking operation (Algorithm 1) confines each pattern $p$ to cluster $C\in\mathcal{C}_{k}$, shrinking candidates from $\lvert D\rvert$ to $\lvert C\rvert$. Accordingly, the decision space contracts from
$O(|I|\,|D|+|P|\,|D|)\longrightarrow\sum_{k}\ \sum_{C\in\mathcal{C}_{k}}\ O\!\big(|I_{C}|\,|C|+|P_{C}|\,|C|\big),$
where $I_{C}$ and $P_{C}$ denote the subsets of items and patterns confined to cluster $C$ (by sinking). Under this clustering, global max-latency constraints are satisfied within each layer, eliminating cross-layer coupling in the optimization.
(ii) Cluster-level separability weakens objective coupling. Within each cluster, WAN bandwidth is far more homogeneous. Hence, the heterogeneity-driven coupling term in $C^{(A)}$ (Section III-A, Equation (5)) is largely attenuated intra-cluster, making the total objective approximately additive across clusters and enabling parallel per-cluster solving.
(iii) Overlap-centric placement lowers granularity. For non-replicated patterns, we further decompose them into disjoint overlap regions. At layer $L_{k}$, let $\mathcal{C}_{k}$ be the set of clusters. For a cluster $C\in\mathcal{C}_{k}$, let $BS_{C}$ denote its candidate bridge subgraphs and $\{\omega\}_{\mathcal{C}}$ the set of disjoint regions. The DHD model provides a score for each region $\omega$, based on which a single-criterion greedy assignment is applied. The per-region cost is
$O\big(|BS_{C}|+T_{\mathrm{DHD}}(C)\big),$
where $T_{\mathrm{DHD}}(C)$ is the time required for the diffusion to reach a steady state on cluster $C$. Let $E_{C}^{B}$ be the set of cross-partition edges in $C$ used by the DHD computation and candidate filtering. The total complexity of the placement phase is thus
$O\!\left(\sum_{k}\ \sum_{C\in\mathcal{C}_{k}}\Big(|E_{C}^{B}|\;+\;|\{\omega\}_{\mathcal{C}}|\cdot\big(|BS_{C}|+T_{\mathrm{DHD}}(C)\big)\Big)\right).$
(iv) Stepwise routing prunes the search tree. Routing is handled in stages. In the online mode, bottom-up greedy selection within each layer quickly prunes participating DCs by maximizing coverage before escalating to upper layers. In the offline mode, routing first performs top-down localization and then bottom-up assembly; a local migration test (Section VI, Equation (14)) replaces an exponential global subset search across all DCs, thereby avoiding combinatorial explosion.
IX-D Guarantees and Decision Rules
This section discusses how the components of $\mathsf{GeoLayer}$ optimize the objective function and presents the associated approximation guarantees.
Since the joint objective in Equation (6) is NP-hard and non-submodular, we do not claim a global constant-factor approximation to Equation (6). Instead, $\mathsf{GeoLayer}$ establishes: (i) feasibility by construction; (ii) monotone non-increase of a principled surrogate objective at each actionable step; and (iii) cluster-level near-additivity, which renders subproblems separable and enables parallel optimization.
Feasibility. Latency requirements of patterns are enforced by layering constraints. Capacity/budget checks are verified before committing replicas or routes. Any action violating a hard constraint is rejected.
Monotone non-increase. For each candidate action $\tau$, we evaluate the marginal change $\Delta\widehat{C}(\tau)=\Delta C^{R}(\tau)-\big(\Delta C^{S}(\tau)+\Delta C^{W}(\tau)+\Delta C^{A}(\tau)\big)$. Following Equation (13) in replica placement and Equation (14) in stepwise routing, we accept $\tau$ only if $\Delta\widehat{C}(\tau)<0$; otherwise $\tau$ is a no-op. Hence every committed action is non-increasing in $\widehat{C}$, and strictly improving whenever the marginal gain is strictly negative.
Cluster-level near-additivity and parallelism. Within a layer, bridge subgraphs induce latency-homogeneous clusters, which attenuate the heterogeneity term in $C^{(A)}$ (Section III-A, Equation (5)). Consequently, the objective is approximately additive across clusters, rendering the subproblems separable and enabling parallel optimization.
Approximation (surrogate-level). Under standard cluster-local assumptions (monotone diminishing-returns read benefit and additive storage/write/association costs), the regional greedy admits the classic (11/e)(1-1/e) approximation to the surrogate objective under a cardinality constraint [90]. We emphasize that this guarantee applies to the surrogate, not to Equation (6) in full generality.
IX-E Implementation
This section presents the full implementation of Overlap-Centric Replica Placement in $\mathsf{GeoLayer}$. We decompose the process into three components: the pattern replica sinking operation (Algorithm 1), the Overlap-Centric Replica Placement (Algorithm 2), and the eviction of cold replicas during online query execution (Algorithm 3).
Algorithm 1 describes the pattern replica sinking process. The input pattern set $P$ is derived from historical access logs and carries request frequency and latency requirement features. The layer set $L$ is constructed using a fixed-threshold bucketing scheme (e.g., 100 ms per interval). If layer $L_{k}$ meets the latency requirements of pattern $p$, then $p$ is placed into all bridge subgraphs in this layer that request $p$ (Line 1). Otherwise, $p$ will sink to layer $L_{k-1}$ and be replicated across all bridge subgraphs that request it. At layer $L_{k-1}$, its latency requirement will then be examined to determine whether it should continue sinking. (Lines 1-1).
After the pattern replica sinking is completed, $\mathsf{GeoLayer}$ performs the overlap-centric replica placement strategy. This strategy quantifies the benefit of data replication and leverages a directed heat diffusion model to make the final placement decisions. Algorithm 2 outlines the replica placement process. Initially, any external pattern $p$ retrieved at a given layer $L_{k}$ is included in the replication procedure if it satisfies $R_{pL_{k}}>W_{pL_{k}}$. The process begins with the layer-by-layer sinking of pattern replicas, guided by latency requirements (Line 2). Next, the data replication gain $C^{rep}_{p}$ is computed to determine whether the pattern should be fully replicated or decomposed in the next lower layer (Lines 22). If $C^{rep}_{py}<0$, the pattern is decomposed across the corresponding BSs (Line 2). For pattern decomposition, patterns are first partitioned into a set of disjoint regions based on their overlapping components (Line 2). The replication gain ${C}^{rep}_{\omega}$ of each region is then evaluated to determine whether the region should be replicated across $BS$s or enter a competitive allocation process (Line 2). For regions subject to competition, each $BS$ executes the DHD model (Line 2), and region ownership is determined based on the resulting heat distribution (Line 2). This process proceeds layer by layer until all data items reach the DCs at layer $L_{0}$ (Line 2). Finally, high-heat regions are pre-cached in the DCs at layer $L_{0}$ (Line 2).
1
Output: Final pattern distribution $\mathbb{P}$
2
3 Function $\textnormal{{ReplicaSink}}()$
4
5 for $k\leftarrow h$ to 11 do
6     while $\mathbb{P}^{(k)}\neq\emptyset$ do
7        $p\leftarrow\mathbb{P}^{(k)}.\text{pop}()$
8        if $\eta_{p}\cdot\Gamma_{\max}\in(t_{k-1},t_{k})$ then
9           $\textnormal{{PlaceReplica}}(p,\mathbb{P}^{(k)},L_{k})$
10
11        else
12             // Latency violation - sinking
13           $\textnormal{{FullReplicate}}(p,L_{k-1})$
14           $\textnormal{{RecordPlacement}}(p,L_{k-1})$
15           $\mathbb{P}^{(k-1)}\leftarrow\mathbb{P}^{(k-1)}\cup\{p\}$
16
17
18
19 return $\mathbb{P}$
Algorithm 1 Pattern Replica Sinking Process